Solve listed initial value problems by using the Laplace Transform:

7.       yll yl 2y = 3 e2t     y(0) = 1, yl(0) = 5
Answer #1

y - y. 24 cze²t you cut y10125 Apply lapláce transformia Lly} - L{4} - ; 2 L4 y} 3L4: e2tę 8²418) - sy(o) - syco) - {S41yol (8 1268-2). E :3 +48 +! L á ? +482-88 +8-2 _ 8-2 st r . 482- 7.8t 1 , 8-2 YA12 482-78th ...18+Ij487242. By partial fractiYis)= 94. - t 8.l t 318+) 3 18-2) (8-2)² taking inuense laplace trains fourni - 1$ 4.033 Y 1) = 641-1644.+ _14(.2)+ y 15-2)2